1. The city of Flowertown (pop: 135) produces and consumes two goods: cereal and milk. It produced these goods in the following amounts during 2015, 2016, and 2017.  

Year Cereal Milk
Quantity Price Quantity Price
2015 200 $1.25 160 $1.20
2016 225 $1.40 160 $1.25
2017 235 $1.45 180 $1.45

If 2015 is the base year then

a.

the Nominal GDP and Real GDP for 2015 would be $192.

b.

the Nominal GDP and Real GDP for 2015 would be 100.

c.

the Nominal GDP and Real GDP for 2015 would be $442.

d.

the Nominal GDP and Real GDP for 2015 would be $250.

2. The city of Flowertown (pop: 135) produces and consumes two goods: cereal and milk. It produced these goods in the following amounts during 2015, 2016, and 2017.  

Year Cereal Milk
Quantity Price Quantity Price
2015 200 $1.25 160 $1.20
2016 225 $1.40 160 $1.25
2017 235 $1.45 180 $1.45

If 2015 is the base year then

a.

the GDP Deflator for 2015 is 442.50 and the GDP Deflator for 2016 is 225.40.

b.

the GDP Deflator for 2015 is 440.32 and GDP Deflator for 2016 is 315.

c.

the GDP Deflator for 2015 and 2016 both are 100.

d.

the GDP Deflator for 2015 is 100 and the GDP Deflator for 2016 is 108.82.

3. Recessions is associated with which of the following?

a.

increase in incomes

b.

decreased bankruptcies

c.

higher unemployment

d.

increasing profits
Answer #1

1) NGDP in

2015 = 200*1.25+160*1.20 = 250+192 = 442

As the base year is 2015 so the Real GDP would also be the same as NGDP = 442

option(C)

2) In the base year, the GDP deflator would always be 100

also, nominal GDP in 2016 = 225*1.4+160*1.25 = 315+200 = 515

Real GDP in 2016 = 225*1.25+160*1.20 = 281.25+192= 473.25

GDP D in 2016 = 515/73.25*100 = 108.82

Option(D)

3) Higher unemployment

option(C)